OpenStudy (anonymous):

If a green ball has a greater momentum than an orange ball and both balls are moving at the same velocity, then.

OpenStudy (zephyr141):

momentum is \[p=mv\] momentum is p. so if one has greater momentum than the other but they are traveling at the same velocity then there's only one other variable that can change.

OpenStudy (michele_laino):

I write the momentum of the green ball as below: \[{p_G} = {m_G}{v_0}\] and the momentum of the orange ball: \[{p_O} = {m_O}{v_0}\] where \(m_G,m_O\) are the masses of the two balls respectively, and \(v_0\) is the common speed. Now I take the ration between the two equation above, and I write: \[\frac{{{p_G}}}{{{p_O}}} = \frac{{{m_G}{v_0}}}{{{m_O}{v_0}}} = \frac{{{m_G}}}{{{m_O}}} > 1\] being, by hypothesis \(p_G>p_O\), so its ratio is greater than \( 1\)